OpenStudy (learn123):

The number of subsets that can be created from the set {1, 2, 3} is: 3 6 7 8

zepdrix (zepdrix):

\[\large\rm S=\{1,2,3\}\] The cardinality of the power set tells us the total number of subsets that can be created.\[\large\rm |\mathcal P(S)|=2^n\]It's 2 raised to the power of the number of elements in your set. We have three elements in our set, so, \(\large\rm 2^3=?\)

OpenStudy (holsteremission):

As for why \(|\mathcal P(S)|=2^n\) if \(|S|=n\), that's because any subset can be made by choosing \(k\) elements from \(S\), and adding up all the ways \(k\) elements can be taken. Suppose \(S=\{1,2,\ldots,n\}\), then you have \[\begin{align*} \text{empty set}&\implies \binom n0\text{ way of choosing a zero-element set}\\ \{1\},\{2\},\ldots,\{n\}&\implies \binom n1\text{ ways of choosing a one-element set}\\ \{1,2\},\{1,3\},\ldots,\{1,n\},\\ \{2,3\},\{2,4\},\ldots,\{2,n\},\ldots,\\ \{n-2,n-1\},\{n-2,n\},\\ \{n-1,n\}&\implies \binom n2\text{ ways of choosing a two-element set}\\ &~~~~\vdots \end{align*}\]and so on. Adding up all the ways means computing the sum \[\sum_{k=0}^n\binom nk=\sum_{k=0}^n\binom nk 1^{n-k}1^k=(1+1)^n=2^n\]where the first second equality relies on the binomial theorem.